. Calib received $50 for his birthday that he put in his piggy bank. Each week, he puts 2 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks. Is this model linear, and, if so, what is the slope and y-intercept?

Answer by haileytucki(390)

If thia model is linear, you need to change the d to x and w to y:
x=2y+50
Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
2y+50=x
Since 50 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 50 from both sides.
2y=-50+x
Move all terms not containing y to the right-hand side of the equation.
2y=x-50
Divide each term in the equation by 2.
(2y)/(2)=(x)/(2)-(50)/(2)
Simplify the left-hand side of the equation by canceling the common factors.
y=(x)/(2)-(50)/(2)
Simplify the right-hand side of the equation by simplifying each term.
y=(x)/(2)-25
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=(1)/(2) and b=-25.
m=(1)/(2), b=-25